OpenStudy (anonymous):

MEDALS. TESTIMONIAL. NEEEEED HELP. ONLY 3. 1.) What is the coordinate of the midpoint of AE? 2.) What is the distance between points A and G? 3.) What is the distance between points C and F?

OpenStudy (anonymous):

midpoint of ae is i

OpenStudy (anonymous):

The answer choices are numbers.

hartnn (hartnn):

then the co-ordinates of all those points should also be numbers

OpenStudy (anonymous):

there arent numbers in ur pic :/

OpenStudy (anonymous):

1: 5, -6, -3, & -1. 2: -10, 10, -17, & 17. 3: 1, -1, 7, & 7. Sorry! I keep putting the wrong picture. Above are the asnwer choices.

hartnn (hartnn):

A is at -10 E is at 0 so, its midpoint will be at the average point, so, whats (-10+0)/2 =.... ?

OpenStudy (anonymous):

-10?! @hartnn

OpenStudy (anonymous):

Is that 12?!

hartnn (hartnn):

no, -10+0 = -10 but -10/2 =... ?

OpenStudy (anonymous):

-5! Wb the next two?!

hartnn (hartnn):

yeah, -5 A = -10 , G = 7 distance = difference = 7- (-10) = .. ?

OpenStudy (anonymous):

17?! @hartnn

OpenStudy (anonymous):

14 i believe

OpenStudy (anonymous):

14 isnt a answer choice?

hartnn (hartnn):

correct! same way, calculate distance between C and F C =-3, F = 4 d= 4- (-3) =... ?

OpenStudy (anonymous):

nope 17 is right lol srry

OpenStudy (anonymous):

7, @hartnn! Thank you!

hartnn (hartnn):

welcome ^_^